# A discrete Benamou-Brenier formulation of Optimal Transport on graphs

### v1

@@ IV Conclusion
[…]
$\mathcal{I}_{q}(v,g)^{q}$ $=\int_{0}^{1}\sum_{k}g(t)_{k}|v(t)_{k}|^{q}dt$
$\geq\left(\int_{0}^{1}\sum_{k}|g(t)_{k}v(t)_{k}|dt\right)^{q}$
$\geq\left(\sum_{k}|\int_{0}^{1}g(t)_{k}v(t)_{k}dt|\right)^{q}$
$=\left(\sum_{k}|P\cdot(f(1)-f(0))+\epsilon|\right)^{q}$
$=\int_{0}^{1}|\tilde{v}|^{q}dt$
$=\int_{0}^{1}\sum_{k\in\mathcal{E}}\tilde{g}(t)_{k}|\tilde{v}(t)_{k}|^{q}dt=\mathcal{I}_{q}(\tilde{v},\tilde{g})^{q}$
@@
So we have additionally that $V_{q}\geq W_{1}$. Conversely, choose $J$ a minimising solution (achieves $W_{1}$) to the Beckmann formulation, then we can choose $|v|=\sum_{k}|J_{k}|$, $v(t)_{k}=\text{sign}(J_{k})|v|$ and $g(t)_{k}=|J_{k}|/|v|$, which satisfies $g(t)_{k}v(t)_{k}=J_{k}$. Additionally this pair $(v,g)$ is constant in time, so we have that
@@
-F Supplementary Results

### v2

@@ V Conclusion
Let $(v,g)$ be a solution to (13), we induce a path $f$, and then since $(f,v,g)$ satisfy the transport equation, for some $(P,\epsilon)$ we have $vg=P\cdot\partial_{t}f+\epsilon$. Then define $(v^{*},g^{*})$ by
$v^{*}_{k}$ $=\text{sign}(P\cdot\partial_{t}f+\epsilon)_{k}|v^{*}|$
$g^{*}_{k}$ $=|P\cdot\partial_{t}f+\epsilon|_{k}/|v^{*}|$
$|v^{*}|$ $=\sum_{k\in\mathcal{E}}|P\cdot\partial_{t}f+\epsilon|_{k}$
then we know $\int_{0}^{1}v^{*}g^{*}dt=P\cdot(f(1)-f(0))+\epsilon$ and so $\int_{0}^{1}\Omega v^{*}g^{*}dt=f(1)-f(0)$. Finally we have
$\mathcal{I}_{q}(v,g)^{q}$ $=\int_{0}^{1}\sum_{k\in\mathcal{E}}g_{k}|v_{k}|^{q}dt$
$\geq\int_{0}^{1}\left(\sum_{k\in\mathcal{E}}|g_{k}v_{k}|\right)^{q}dt$
$=\int_{0}^{1}\left(\sum_{k}|P\cdot\partial_{t}f+\epsilon|_{k}\right)^{q}dt$
$=\int_{0}^{1}|v^{*}|^{q}dt$
$=\int_{0}^{1}\sum_{k}g^{*}_{k}|v_{k}^{*}|^{q}dt=\mathcal{I}_{q}(v^{*},g^{*})^{q}.$
Now consider an increasing function $\phi:[0,1]\rightarrow[0,1]$ such that $\phi(0)=0$ and $\phi(1)=1$ and the following pair:
$v^{**}(t)_{k}=\phi^{\prime}(t)\cdot v^{*}(\phi(t))_{k},\quad g^{**}(t)_{k}=g^{*}(\phi(t))_{k}.$
By integration by substitution we know
$\int_{0}^{1}\Omega\cdot v^{**}g^{**}dt=\int_{0}^{1}\Omega\ \cdot v^{*}g^{*}dt=f(1)-f(0).$
So $(v^{**},g^{**})$ is a valid solution. We consider when
$\left|\phi^{\prime}(t)\frac{dv^{*}}{d\phi}\right|=\phi^{\prime}(t)\left|\frac{dv^{*}}{d\phi}\right|=c.$ (16)
We denote $s(\phi)=\left|\frac{dv^{*}}{d\phi}\right|$ and $S(y)=\int_{0}^{y}s(y)dy$ then solving (16) we have $S(\phi(t))=ct+d$. We know that $c=S(\phi(1))-S(\phi(0))=S(1)-S(0)$ and $d=S(\phi(0))=S(0)$. So the choice of
$\phi(t)\coloneqq S^{-1}[(S(1)-S(0))t+S(0)]$
gives us $|v^{**}|=c=S(1)-S(0)=\int_{0}^{1}s(t)dt$. Note that $S^{-1}$ exists because $s(y)>0$ so $S$ is increasing. This parametrisation has time-independent speed and by the chain rule and Jensen’s inequality we have
$\mathcal{I}_{q}(v^{**},g^{**})^{q}$ $=\int_{0}^{1}|v^{**}|^{q}dt$
$=|v^{**}|^{q}$
$=\left(\int_{0}^{1}|v^{**}|dt\right)^{q}$
$=\left(\int_{0}^{1}\phi^{\prime}(t)|v^{*}(\phi(t))|dt\right)^{q}$
$=\left(\int_{0}^{1}|v^{*}(s)|ds\right)^{q}$
$\leq\int_{0}^{1}|v^{*}(s)|^{q}ds=\mathcal{I}_{q}(v^{*},g^{*})^{q}$
@@
So we have that $V_{q}\geq W_{1}$. Conversely, choose $J$ a minimising solution (achieves $W_{1}$) to the Beckmann formulation, then we can choose $|v|=\sum_{k}|J_{k}|$, $v(t)_{k}=\text{sign}(J_{k})|v|$ and $g(t)_{k}=|J_{k}|/|v|$, which satisfies $g(t)_{k}v(t)_{k}=J_{k}$. Additionally this pair $(v,g)$ is constant in time, so we have that
@@
-F Proof of Local Minimisation
Proof.
Suppose $(v,g)$ is a minimising pair for $I_{1}$ across $[0,1]$, i.e we have that
$\mathcal{I}_{1}(v,g)=\int_{0}^{1}\sum_{k\in\mathcal{E}}v_{k}g_{k}dt$ $=\min_{(v_{0},g_{0})}\mathcal{I}_{1}(v_{0},g_{0})$
$=W_{1}(f(0),f(1)).$
Then we denote
$\mathcal{I}_{q}^{[s,t]}(v,g)=\left(\int_{s}^{t}\sum_{k\in\mathcal{E}}g(t)_{k}|v(t)|_{k}^{q}dt\right)^{\frac{1}{q}}.$
This implies that $\mathcal{I}_{1}^{[0,1]}=\mathcal{I}_{1}$. Suppose $(v,g)$ is not minimal across some $[s,t]$ for $\mathcal{I}_{1}$, then there exists some $(v^{*},g^{*})$ such that
$\mathcal{I}_{1}^{[s,t]}(v^{*},g^{*})<\mathcal{I}_{1}^{[s,t]}(v,g).$
From here we construct a new solution across $[0,1]$:
$(v^{**}(\tau)_{k},g^{**}(\tau)_{k})=\begin{cases}(v^{*}(\tau)_{k},g^{*}(\tau)_{k})&:\tau\in[s,t]\\ (v(\tau)_{k},g(\tau)_{k})&:\tau\not\in[s,t]\end{cases}$
Then by decomposing
$\mathcal{I}_{1}^{[0,1]}=\mathcal{I}_{1}^{[0,s]}+\mathcal{I}_{1}^{[s,t]}+\mathcal{I}_{1}^{[t,1]}$
we have that
$\mathcal{I}_{1}^{[0,1]}(v^{**},g^{**})$ $<\mathcal{I}_{1}^{[0,s]}(v,g)+\mathcal{I}_{1}^{[s,t]}(v,g)+\mathcal{I}^{[t,1]}(v,g)$
$=\mathcal{I}_{1}^{[0,1]}(v,g)$
which contradicts the minimality of $(v,g)$. ∎
-G Supplementary Results